OpenStudy (anonymous):

Select an ordered pair from the choices below that is a solution to the following system of inequalities: 2x + 4y > 8 y > 2x + 3 A. (-2, -3) B. No solution C. (-1, 2) D. (2, 9)

OpenStudy (anonymous):

option D as it satisfies the given set of inequalities

OpenStudy (anonymous):

can you explain how you solve these type of questions please ?

OpenStudy (anonymous):

sketch both the curves first

OpenStudy (anonymous):

taking 2x+4y=8 and y=2x+3

OpenStudy (anonymous):

i agree with you nitz but we need ti check the signs that is. > and < symlos properly

OpenStudy (anonymous):

obviously check out the portion at right common to both the curves]

OpenStudy (anonymous):

one simple way is substitute these x and y values in the equation and check the correct ness of > and < symbols i.e are they satisfying them or not

OpenStudy (anonymous):

i have no clue what that means. can you show me a formula to solve it ?

OpenStudy (anonymous):

i have to trace the graph

OpenStudy (anonymous):

like someone said put the graphs and check the point of intersections such that it forms an interval and verify if these ordered pairs are contained in these intervals

OpenStudy (anonymous):

there are many ways u can even solve it by equations

OpenStudy (anonymous):

if u need to approach in a conceptual way better go with graphical representation

OpenStudy (anonymous):

there has to be a formula because they wouldnt give me a question with no graph

OpenStudy (anonymous):

are you new to this inequalities stuff?? b'coz it needs a lot of practice and there are several rules to solve such problems

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

you need to draw graphs on your own..none of them is gonna provide you..typical and greatest problems of maths are solve by using graphs

OpenStudy (anonymous):

ok but what am i supposed to be graphing ? & how ?

OpenStudy (anonymous):

you should be graphing the given equations

OpenStudy (anonymous):

the equation here says 2x+4y>8 first graph 2x+4y=8 then go for values which are greater than this curve and likewise

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

so is the problem here graphing? or still finding the ordered pair?

OpenStudy (anonymous):

 here it is the graph for 2x+4y=8 the arrows pointing upward means 2x+4y>8 and likewise draw for the other equation in the same graph. check the point of intersections and evaluate the results by considering the ordered pairs which you have provided as the options